OpenStudy (anonymous):

How do I simplify √1/4x^2?

OpenStudy (unklerhaukus):

is that \[\sqrt{\frac1{4x^2}}\]?

OpenStudy (unklerhaukus):

\[\sqrt{\frac1{4x^2}}=\frac{\sqrt1}{\sqrt{4x^2}}=\frac{\sqrt1}{\sqrt{4}\sqrt{x^2}}= \dots\]

OpenStudy (anonymous):

Oh so would the final answer be x/2 ?

OpenStudy (unklerhaukus):

almost,

OpenStudy (anonymous):

SO you multiply 2x on both sides, giving us 2x? or do you have to do something with the √1?

OpenStudy (unklerhaukus):

don't multiply anything, just simplify each square root term

OpenStudy (anonymous):

√1/2x

OpenStudy (unklerhaukus):

yeah that's right, and you can simplify √1

OpenStudy (anonymous):

The only thing I don't understand is that the x does not have a specific place, so can we determine that it is the bottom or the top, and if so how?

OpenStudy (unklerhaukus):

oh, i interpreted your question wrong.

OpenStudy (unklerhaukus):

\[\sqrt{\frac14x^2}=\sqrt{\frac14}\sqrt{x^2}=\frac{\sqrt1}{\sqrt4}\sqrt{x^2}=\frac{1}{2}x=x/2\] as you said originally

OpenStudy (anonymous):

Ok, thank you SO much!

OpenStudy (unklerhaukus):

NB we have only computed the principal square roots